# ODGEN: Domain-specific Object Detection Data Generation with Diffusion Models

### v1

@@ D.2 ODGEN (ours)
We fine-tune the parameters in the UNet of Stable Diffusion with the default training setting used by ControlNet [60], including the learning rates and the optimizer.
@@
Image encoder. ControlNet uses a 4-layer convolutional encoder to encode input conditions sized $512$ to $64\times 64$ latents. The input channels are usually set as 1 or 3 for input images. Our ODGEN follows similar methods to convert the input conditions. The key difference is that we use image lists of synthesized objects as input conditions. The image lists are padded to $N$ and concatenated in the channel dimension for encoding. In this case, we expand the input channels of the encoder to $3N$ and update the channels of the following layers to comparable numbers. We provide the channels of each layer in Tab. 9. The convolutional kernel size is set as 3 for all layers.
@@ D.3 Other Methods
ReCo [59] & MIGC [65] & GeoDiffusion [5] are layout-to-image generation methods based on Stable Diffusion. For the comparison on RF7, we use their official codes to fine-tune the diffusion models on domain-specific datasets and generate synthetic datasets for experiments. For the comparison on COCO, we directly use their released pre-trained weights trained on COCO. MIGC only releases the pre-trained weight but not the code, so we only compare it with our method on the COCO benchmark. ReCo and GeoDiffusion fine-tune both the text encoder and the UNet of Stable Diffusion. MIGC only fine-tunes the UNet.
@@
     $left,top,width,height=bbox[i]$
@@
     $M[top:top+height,left:left+width,k]+=1$
@@ Appendix E Datasets Details
Appendix F Supplemental Experiments
Other related works. InstaGEN [11] is mainly designed for the open-vocabulary object detection task. It generates images randomly and annotates them with a grounding head trained with another pre-trained detector. InstanceDiffusion [53] needs detailed segmentation masks for training and inference. They are not included for direct comparison in our paper.
@@ Appendix F Supplemental Experiments
Visualization. We provide extensive visualized examples of ODGEN to demonstrate its effectiveness. In Fig. 9, we show generated samples of the specific domains in RF7. It can be seen that ODGEN is robust to complex scenes composed of multiple categories, dense and overlapping objects. Besides, ODGEN shows strong generalization capability to various domains.

### v2

@@ D.2 ODGEN (ours)
We follow the default training setting used by ControlNet [62], including the learning rates and the optimizer to train our ODGEN models and provide fair comparison with ControlNet.
@@
Image encoder. ControlNet uses a 4-layer convolutional encoder to encode input conditions sized $512$ to $64\times 64$ latents. The input channels are usually set as 1 or 3 for input images. Our ODGEN follows similar methods to convert the input conditions. The key difference is that we use image lists of synthesized objects as input conditions. The image lists are padded to $N$ and concatenated in the channel dimension for encoding. In this case, we expand the input channels of the encoder to $3N$ and update the channels of the following layers to comparable numbers. We provide the channels of each layer in Tab. 17. The convolutional kernel size is set as 3 for all layers. We design different encoder architectures according to the maximum object numbers that can be found in a single image. For datasets like MRI in which most images contain only one object, we can use a smaller encoder to make the model more lightweight. For a generic model trained on large-scale datasets, we can fix the architectures of encoders with a large $N$ to make it applicable to images containing many objects.
@@ D.3 Other Methods
ReCo [61] & MIGC [67] & GeoDiffusion [5] are layout-to-image generation methods based on Stable Diffusion. For the comparison on RF7, we use the official codes of ReCo and GeoDiffusion to fine-tune the diffusion models on domain-specific datasets and generate synthetic datasets for experiments. For the comparison on COCO, we directly use their released pre-trained weights trained on COCO. MIGC only releases the pre-trained weight but not the code, so we only compare it with our method on the COCO benchmark. ReCo and GeoDiffusion fine-tune both the text encoder and the UNet of Stable Diffusion. MIGC only fine-tunes the UNet.
InstanceDiffusion [55] depends on multiple formats of inputs, including segmentation masks, boxes, and scribbles, to train its UniFusion module. However, the RF7 datasets of specific domains do not provide segmentation masks and scribbles. Therefore, it is only employed for comparison on COCO with its open-source checkpoint.
@@
     $coord_{left},coord_{top},coord_{right},coord_{bottom}=bbox[i]$
@@
     $M[coord_{top}:coord_{bottom},coord_{left}:coord_{right},k]+=1$
@@ Appendix E Datasets Details
Figure 10: Left: images used to build image lists. Right: corresponding synthesized samples. The synthesized images of foreground objects share only the same category with objects in synthesized samples but not colors or shapes.
Figure 11: Detection results of YOLO detectors trained w/ or w/o ODGEN synthetic data. The synthesized samples of ODGEN help detectors perform better on partially occluded objects.
Figure 12: Visualized samples containing novel categories of foreground objects generated by ODGEN trained on the COCO dataset. Stable Diffusion is used to generate images of the foreground objects to build image lists for inference. It shows that our ODGEN can control the layout of novel categories that were never seen in its training process.
Figure 13: Visualized samples generated by the models fine-tuned on both cropped foreground objects and entire images from RF7 datasets (200 images for each dataset). Text prompts are provided under figures. Columns 1-6 are generated foreground objects and columns 7-9 are generated entire images. Models fine-tuned on entire images only cannot synthesize foreground objects required by image lists since they cannot bind the prompts to corresponding objects in entire images that may contain multiple categories of objects.
@@
Appendix F Supplemental Related Works
Copy-paste [14, 65] method requires segmentation masks to get the cropped foreground objects, which are not provided by the RF7 datasets used in this paper. Besides, our approach also serves as a controllable image generation method that only needs bounding box labels to produce realistic images, which is hard to achieve by copy-paste. InstaGEN [11] is mainly designed for the open-vocabulary object detection task. It generates images randomly and annotates them with a grounding head trained with another pre-trained detector. Earlier works like LayoutDiffusion [66] are not implemented with latent diffusion models and thus are not included for comparison. Works [23, 39, 56] are designed for semantic segmentation, while our ODGEN is designed for object detection.
Appendix G Supplemental Experiments
@@
Visual relationship between image lists and synthesized images. The image list is designed to provide information on category and localization for controlling foreground object layouts. The synthesized images share the same category with objects in generated images and are pasted to the same position as those in generated images to build image lists. The objects in image lists and synthesized images may all be apples or cakes but have different shapes and colors, as shown by the visualized samples provided in Fig. 11.
Detector performance improvement with synthetic data. We provide several visualized samples in Fig. 11 and show that the synthetic data helps detectors detect some occluded objects.
Generalization to novel categories. We use ODGEN trained on COCO to synthesize samples containing subjects not included in COCO, like moon, ambulance, tiger, et al. We show visualized samples in Fig. 12. It shows that ODGEN trained on COCO can control the layout of novel categories. However, we find that the generation quality is not very stable, which may be influenced by the fine-tuning process on the COCO dataset. In future work, we hope to get a powerful and generic model that is robust to the most common categories in daily life.
Computational cost. Our ODGEN needs to synthesize images of foreground objects to build image lists, leading to higher computational costs for training and inference. Therefore, we propose to generate an offline image library of foreground objects to accelerate the process of building image lists. With the offline library, we can randomly pick images from it to build image lists instead of synthesizing new images of foreground objects every time.
Taking the model trained on the COCO dataset as an example, our ODGEN shares a very close scale of parameters with ControlNet (Trainable parameters: ODGEN 1231M v.s. ControlNet 1229M, parameters in the UNet of Stable Diffusion are included). We provide the training time for 1 epoch on COCO with 8 V100 GPUs of different methods in Tab. 21.
Table 21: Training time for 1 epoch on COCO with 8 V100 GPUs.
Method ReCo GLIGEN ControlNet GeoDiffusion ODGEN
Training Time 3 hours 7 hours 4.5 hours 3.2 hours 5.6 hours
In the inference stage (Fig. 2(c)), our ODGEN pads both the image and text lists to a fixed length. Therefore, the computational cost for inference with an offline library doesn’t increase significantly with more foreground objects. Other methods like InstanceDiffusion [55] and MIGC [67] need more time for training and inferencing with more objects. Taking models trained on COCO as examples, to generate an image with 6 bounding boxes on a V100 GPU, ODGEN takes 10 seconds, ControlNet takes 8 seconds, and InstanceDiffusion takes 30 seconds.
During the inference process, we compare using the same offline image library as training against using a totally different image library. We get very close results, as shown in Tab. 22. Both outperform other methods significantly, as shown in Tab. 4. It indicates that ODGEN can extract category information from synthesized samples of foreground objects in image lists instead of depending on certain images of foreground objects. ODGEN is not constrained by the image library of foreground objects used in training and can be generalized to other newly generated offline image libraries consisting of novel synthesized samples of foreground objects, which ensures that offline image libraries increase ODGEN’s computational efficiency without reducing its practicality.
Table 22: FID ($\downarrow$) and mAP ($\uparrow$) of YOLOv5s / YOLOv7 for ODGEN trained on COCO. FID is computed with 41k synthetic images. For mAP, YOLO models are trained from scratch on 10k synthetic images and validated on 31k real images. Different offline image libraries are applied for inference.
Offline Image Library FID ($\downarrow$) mAP@.50 ($\uparrow$) mAP@.50:95 ($\uparrow$)
Same as training 16.01 18.90 / 9.70 24.40 / 14.20
Different from training 16.16 18.60 / 9.52 24.20 / 14.10
Quantitative evaluation for concept bleeding. This work mainly focuses on the concept of the bleeding problem of different categories of objects. The mAP results of YOLO models trained on synthetic images only in Tab. 4 can be a proxy task to prove that the concept bleeding is alleviated since our ODGEN achieves state-of-the-art performance in the layout-image consistency of complex scene generation conditioned on bounding boxes. This section adds quantitative evaluation with BLIP-VQA [21], which employs the BLIP [32] model to identify whether the contents in synthesized images are consistent with text prompts. The results are provided in Tab. 23. Our ODGEN outperforms other methods and gets results close to ground truth (real images from the COCO validation set sharing the same labels with synthetic images).
Table 23: BLIP-VQA ($\uparrow$) results averaged over 41k images synthesized with different methods using the labels from the COCO validation set.
Method ReCo GLIGEN ControlNet GeoDiffusion
BLIP-VQA ($\uparrow$) 0.2027 0.2281 0.2461 0.2114
Method MIGC InstanceDiffusion ODGEN Ground Truth (reference)
BLIP-VQA ($\uparrow$) 0.2314 0.2293 0.2716 0.2745
ODGEN trained on COCO with Stable Diffusion v1.5. Our ODGEN is implemented with Stable Diffusion v2.1 in Sec. 4. We add experiments of implementing ODGEN with Stable Diffusion v1.5 trained on COCO for comparison in this part. Results are provided in Tab. 24. Our ODGEN achieves similar results with different versions of Stable Diffusion models.
Table 24: FID ($\downarrow$) and mAP ($\uparrow$) of YOLOv5s / YOLOv7 on COCO. FID is computed with 41k synthetic images. For mAP, YOLO models are trained from scratch on 10k synthetic images and validated on 31k real images.
Method FID ($\downarrow$) mAP@.50 ($\uparrow$) mAP@.50:95 ($\uparrow$)
ODGEN based on Stable Diffusion v2.1 16.16 18.90 / 24.40 9.70 / 14.20
ODGEN based on Stable Diffusion v1.5 15.93 18.20 / 24.50 9.39 / 14.30
@@
Visualization. We provide extensive visualized examples of ODGEN to demonstrate its effectiveness. In Fig. 13, we provide visualized samples produced by models fine-tuned on both cropped foreground objects and entire images. It shows that the fine-tuned models are capable of generating diverse foreground objects and complete scenes. In Fig. 14, we show generated samples of the specific domains in RF7. It can be seen that ODGEN is robust to complex scenes composed of multiple categories, dense and overlapping objects. Besides, ODGEN shows strong generalization capability to various domains.